# From Light-Cone to Supersonic Propagation of Correlations by Competing Short- and Long-Range Couplings

### v1

@@ Abstract
Figure 3: (a) The space-time propagation of the correlations $\mathcal{C}_{nn}(d,t)$, in the presence of dissipation, Eqs. (1)-(2), for $\hbar\Omega/J=0.99$, $\hbar\Gamma/J=10$, $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$. The blue line is a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$ and $\Gamma=0$. (b) The propagation of correlations when cavity field is replaced with a classical field supp , for the same parameters as in panel (a). Correlations calculated from TWA as function of time and distance in (c) 1D for a lattice with L=21 and in (d) 2D for lattice with $L\times L=21\times 21$. For the TWA simulations we have used $U/J=0.1$, $\hbar\delta/J=2$, $\hbar\Gamma/J=1$, $\sqrt{N}\hbar\Omega/J=4$ in (c), and $\sqrt{N}\hbar\Omega/J=8$ in (d). For the simulations we initialize the atoms with alternating densities of $n=10$ and $n=11$ bosons corresponding to the total atom numbers (c) $N=220$ and (d) $N=4630$.
@@
Influence of a dissipative cavity field: A seldom explored question is how the presence of dissipation alters the spreading of correlations. Seminal works showed that for local and quasi-local Lindblad operators Lieb-Robinson-like bounds exist Poulin (2010); Schuch et al. (2011); Barthel and Kliesch (2012); del Campo et al. (2013); Funo et al. (2019), observed these numerically Bernier et al. (2018) and characterized the dynamics of mutual information Alba and Carollo (2021) and operator entanglement Wellnitz et al. (2022). The dissipation range has been shown to be able to alter the spatio-temporal behavior of correlations in spin systems Seetharam et al. (2022). Thus, it is an interesting question to understand the effects of dissipation in form of cavity losses, Eq. (1), which is a highly non-local dissipation for the atoms due to the global coupling.
@@
In the present work, the dissipation can be a key player in changing the position of crossover between the local and the global spreading of correlations. By increasing the strength of dissipation we can go from the regime in which the distance-independent propagation of correlations dominates to a regime with a mostly light-cone evolution [see Fig. 3(a) compared to Fig. 1(e)-(f)]. Thus, whereas the dissipation globally couples to the atoms, the photon losses rescale the effective global atom-atom interactions to a lower value [see the coefficient in $\hat{H}_{\text{glo}}$ Eq. (4)], making the light-cone propagation more prominent. However, for the same strength of $\hat{H}_{\text{glo}}$ stronger dissipation helps the supersonic propagation.
@@
Key ingredients for the supersonic propagation: To highlight the required ingredients for competition between the light-cone and distance-independent spreading of the correlations, we show in Fig. 3(b) results obtained employing a mean-field description of the cavity field. Within this approach the cavity mode is modeled by a time-dependent classical field coupled to the mean value $\langle\hat{\Delta}\rangle$ supp . We observe that in this simplified approach we cannot recover the evolution shown in Fig. 3(a). Whereas the light-cone like evolution is evident also in Fig. 3(b), the distance-independent spreading of correlations is absent. As the cavity field couples solely to the mean value $\langle\hat{\Delta}\rangle$ it looses the ability to transport and create fluctuations. However, by including stochastic noise arising from photon losses in the evolution of the classical field we recover a distance-independent propagation, albeit with a reduced contribution than in Fig. 3(a) supp . Thus, the fluctuations of the cavity field are of crucial importance.
Spreading of correlations at higher fillings and two-dimensions: In order to obtain a simplified description which captures the ability of the cavity field to transport density fluctuations, we employ the truncated Wigner approximation (TWA). In the TWA the cavity field and each atomic site is described as a complex stochastic classical field Gardiner et al. (2002); supp and it has been previously shown to describe light-cone spreading of correlations in Bose-Hubbard models Nagao et al. (2019). This approach allows us to extend our tMPS results to regimes of high atomic occupancies and to higher dimensions.
@@
For the 1D case at high filling, we show in Fig. 3(c) the TWA results, for an initial state with a staggered atomic occupation alternating between $n\!=\!10$ and $n\!=\!11$ supp . We highlight that the qualitative behavior is the same as discussed in Fig. 1(e) for the tMPS results, at short times light-cone spreading of the correlations occurs, followed by a distance-independent rise of correlations at large distances. Establishing the TWA as a valid approach to simulate the quantum dynamics of correlations in this model gives us the opportunity to study this effect in higher dimensions. In Fig. 3(d), we show the results for a 2D atomic system coupled to the cavity field, where the atomic initial state is a checkerboard with alternating $n\!=\!10$ and $n\!=\!11$ occupations. Compared to the 1D simulations the magnitude of $\mathcal{C}_{nn}$ is reduced, which we attribute to radial spread of the correlations in 2D. Nevertheless, the key features that emerge from the competing short- and long-range processes are well visible in 2D, the light-cone propagation of correlations on short times and the supersonic spreading on longer times.
@@
Conclusions: We investigated the dynamics of density-density correlations emerging from the interplay of short- and global-range couplings, finding two key features, a light-cone propagation at short distances and times followed by a supersonic, distance-independent, spreading. The crossover time between these dynamical features and their relative strength depend crucially on the ratio between the global coupling and the short-range tunneling. Furthermore, the general character of our results is highlighted by the robustness to the different quench situations, the dimensionality of the system, the presence of dissipation, or the form of the global coupling. We identify that the emerging dynamics relies on the presence of fluctuations in the long-range couplings, stemming either from the cavity field or the interaction itself. The role of the fluctuations is essential for propagating the correlations outside of the light-cone, which further underline the importance of capturing the fluctuations in the atoms-cavity coupling in obtaining the correct quantum dynamics, as shown also in our previous works Halati et al. (2020a, 2022, 2025). The effect could be realized and investigated in current state-of-the-art experiments of ultra-cold atoms coupled to optical cavities. However, due to its generality our findings are applicable to a much wider class of experimental systems. We emphasize that our results cannot be inferred from considerations based on general bounds on the propagation of correlations, which would be dominated by the contributions of the global-range couplings, such that numerical simulations of the quantum dynamics were needed. Our results can guide the experimental investigations in controlling the spreading of correlations in platforms with competing interaction scales.
@@
Acknowledgments: We thank J.P. Brantut and T. Donner for fruitful discussions. We acknowledge support by the Swiss National Science Foundation under Division II Grant No. 200020-219400 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277625399-TRR 185 OSCAR (“Open System Control of Atomic and Photonic Matter”, B4), No. 277146847-CRC 1238 (“Control and dynamics of quantum materials”, C05), CRC 1639 NuMeriQS (“Numerical Methods for Dynamics and Structure Formation in Quantum Systems”) – project No. 511713970 , Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”), and under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769. This research was supported in part by the National Science Foundation under Grants No. NSF PHY-1748958 and PHY-2309135.
@@ A Time-dependent matrix product states method for long-range quantum systems
In order to ensure the convergence of the density-density correlations, $\mathcal{C}_{nn}$, [see main text Eq. (3) for their definition], up to an error of approximately ${}10^{-5}$ for the times considered in the results presented in the absence of dissipation, we chose the following convergence parameters: a maximal bond dimension of $400$ states, which ensured a truncation error of at most $5\times 10^{-8}$ at the final time, a time-step of $dtJ/\hbar=2.5\times 10^{-3}$, the local Hilbert space of the bosonic atoms is $N_{\text{bos}}=4$, and the adaptive cutoff of the local Hilbert space of the photonic mode ranged between $N_{\text{pho}}=70$ and $N_{\text{pho}}=8$. For the parameters used in the presence of dissipation a truncation error of $10^{-8}$ at the final time was achieved for the mentioned convergence parameters, the presented results are averaged over at least 750 quantum trajectories.
The numerical results for the time-evolution of the atomic model with global range interactions, Eq. (4) in the main text, were obtained with an implementation based on matrix product states of the two-site version of the time-dependent variational principle approach (TDVP) Haegeman et al. (2011, 2016), within the ITensor Library Fishman et al. (2022). The convergence was assured by the following parameters of the method: a maximal bond dimension of $250$ states, which ensured a truncation error of at most $10^{-7}$ at the final time, a time-step of $dtJ/\hbar=5\times 10^{-3}$, and the local Hilbert space of the bosonic atoms are $N_{\text{bos}}=4$.
@@ 1 In the limit of small coupling
[…]
@@
×$\qquad\times e^{-t^{2}/2\hbar[\hat{H}_{\text{kin}}+\hat{H}_{\text{int}},\hat{H}_{\text{ac}}]}\dots,$
@@
×$~{}~{}\times\left[1+it\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta}\right]e^{-i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}\big{\rangle}+O(t^{2}\Omega)$
@@
2 Simplified model for the description of the short-time density-density correlation growth
In this section, we aim to capture the minimal ingredients necessary for understanding the distance-independent spreading of the correlations and derive the short-time scaling for the simplified model, $\mathcal{C}_{nn}(d,t)\propto J^{4}\Omega^{4}t^{8}$. As discussed in the main text, when starting from an uncorrelated Fock state the kinetic processes are responsible for creating density-density correlations locally and the coupling to the cavity determines the supersonic spreading. Thus, as a first approximation we neglect the on-site repulsive interactions and the detuning $\delta$ to explain the short time behaviour. In this situation, the 1D version of the Hamiltonian in Eq. (2) of the main text can be written in the momentum basis as
@@ 2 Simplified model for the description of the short-time density-density correlation growth
$\cos(\hat{\theta})=$ $\frac{J+\hat{E}}{\sqrt{2\hat{E}(\hat{E}+J)}},$ (B.14)
$\sin(\hat{\theta})=$ $\frac{\hat{\Omega}}{\sqrt{2\hat{E}(\hat{E}+J)}}.$ (B.15)
@@
where we used Eqs. (B.12) and (B.13) and ${\sin(2\hat{\theta})=\hbar\hat{\Omega}/\hat{E}}$. Consequently, we can calculate for two even sites $l,m$ the density-density correlations with
@@
$\mathrm{Var}(\hat{C})$ $\approx\frac{J^{4}\Omega^{4}t^{8}}{9\hbar^{2}}\mathrm{Var}([\hat{a}+\hat{a}^{\dagger}]^{2}),$ (B.21)
@@ C Treating the cavity mode as a classical field
The comparison between the exact results and the classical field approach is also shown in Fig. 3. As discussed in the main text, by the coupling to the cavity field for the parameters shown the dynamics of the correlations is dominated by a supersonic propagation [Fig. 3(a) and Fig. 3(b)], while for the classical field approach we only observe a light-cone spreading of the correlations [Fig. 3(c) and Fig. 3(d)].
@@
where $\xi(t)$ is a random complex number of magnitude $1$ sampled from a uniform distribution at each point in time. For the initial state $\langle a\rangle(0)$ we use a uniformly sampled random complex number. We simulate different realizations of the time-dependent noise term, similar to the quantum trajectories approach, and average over the values of the computed observables for the different realizations. We obtain that in the presence of the stochastic noise, Fig. C1(e) and Fig. C1(f) it exhibits for larger distances features of the distance-independent rise of the correlations, as seen in Fig. C1(f). This can be understood by the fact that the noise term appearing in Eq. (C.3) translates to a correlated noise term in the evolution of the atoms acting on all sites. However, the contribution from the distance-independent rise of the correlations is much smaller for the shown parameters compared to the full quantum dynamics. This implies that the noise term is insufficient to capture the full dynamics which requires, in particular, also the cavity-mediated transport of density fluctuations.
D Truncated Wigner approximation

### v2

@@ Abstract
Figure 3: (a) The propagation of $\mathcal{C}_{nn}(d,t)$, in the presence of dissipation, Eqs. (1)-(B.1), for $\hbar\Omega/J=0.99$, $\hbar\Gamma/J=10$, $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$. The blue line is a guide to the eye, approximating the front of the light-cone propagation for $\Omega=0$ and $\Gamma=0$. (b) The propagation of correlations when cavity field is replaced with a classical field supp , for the same parameters as in (a). Correlations $\mathcal{C}_{nn}(d,t)$ calculated from TWA in (c) 1D for a lattice with L=21 and in (d) 2D for lattice with $L\times L=21\times 21$. For the TWA simulations we have used $U/J=0.1$, $\hbar\delta/J=2$, $\hbar\Gamma/J=1$, $\sqrt{N}\hbar\Omega/J=4$ in (c), and $\sqrt{N}\hbar\Omega/J=8$ in (d). We initialize the atoms with alternating densities of $n=10$ and $n=11$ bosons corresponding to the total atom numbers (c) $N=220$ and (d) $N=4630$.
@@
Influence of a dissipative cavity field: A seldom explored question is how the presence of dissipation alters the spreading of correlations. Seminal works showed that for quasi-local Lindblad operators Lieb-Robinson-like bounds exist Poulin (2010); Schuch et al. (2011); Barthel and Kliesch (2012); del Campo et al. (2013); Funo et al. (2019), observed these bounds numerically Bernier et al. (2018) and characterized entanglement measures Alba and Carollo (2021); Wellnitz et al. (2022). The dissipation range can further alter the spatio-temporal behavior of correlations Seetharam et al. (2022). Thus, we aim to understand the effects of dissipation in form of cavity losses, Eq. (1), which is a highly non-local dissipation for the atoms due to the global coupling.
@@
For the atoms-cavity system, Eqs. (1)-(B.1), photonic dissipation alters the position of the crossover between the local and the global spreading of correlations. By increasing the strength of dissipation we go from the regime in which the supersonic propagation dominates to a regime with a mostly light-cone evolution [see Fig. 3(a) compared to Fig. 1(e)-(f)]. Thus, whereas the dissipation globally couples to the atoms, the photon losses rescale the effective global atom-atom interactions to a lower value [see the coefficient in $\hat{H}_{\text{glo}}$ Eq. (5)], making the light-cone propagation more prominent. However, for the same strength of $\hat{H}_{\text{glo}}$ stronger dissipation helps the supersonic propagation.
@@
Key ingredients for the supersonic propagation: Our results show the presence of the supersonic propagation for global coupling to the cavity and atom-only global interactions. Thus, we aim to identify the underlying carrier of the correlations, using different approximations to highlight the required ingredients. In Fig. 3(b) we show results obtained employing a mean-field description of the cavity field, for which the cavity mode is modeled by a time-dependent classical field coupled to the mean value $\langle\hat{\Delta}\rangle$ supp . In this mean-field approach we can only recover the light-cone evolution while the distance-independent spreading of correlations is absent [see Fig. 3(b)]. This is due to the inability of the cavity in the mean-field description to transport and create fluctuations. Note that also in the minimal model, Eq. (4), the amplitude of the supersonic spreading would vanish in the mean-field approximation. Thus, we identify fluctuations, which can have different origins, as the carriers of the supersonic correlations. One source of fluctuations in the cavity field is the stochastic noise arising from photon losses. By including such a term in the mean-field description, which translates to a correlated noise in the atomic evolution, we recover a distance-independent propagation, albeit with a reduced contribution than in Fig. 3(a) supp .
To analyze the role of the fluctuations arising in the atoms-cavity coupling we employ the truncated Wigner approximation (TWA). In the TWA the cavity field and each atomic site is described as a complex stochastic classical field Gardiner et al. (2002); supp and it has been previously shown to describe light-cone spreading of correlations in Bose-Hubbard models Nagao et al. (2019). For the 1D case at high filling, we show in Fig. 3(c) the TWA results, for an initial state with a staggered atomic occupation alternating between n=10n\!=\!10 and n=11n\!=\!11 supp . We highlight that the qualitative behavior is the same as discussed in Fig. 1(e) for the tMPS results: at short times light-cone spreading of the correlations occurs, followed by a distance-independent rise of correlations at large distances. Thus, the fluctuations of the global coupling through the cavity field determine the supersonic spreading.
@@
Spreading of correlations in two-dimensions: Establishing the TWA as a valid approach to simulate the quantum dynamics of correlations in this model gives us the opportunity to study this effect in higher dimensions. In Fig. 3(d), we show the results for a 2D atomic system coupled to the cavity field, where the atomic initial state is a checkerboard with alternating n=10n\!=\!10 and n=11n\!=\!11 occupations. Compared to the 1D simulations the magnitude of $\mathcal{C}_{nn}$ is reduced, which we attribute to radial spread of the correlations in 2D. Nevertheless, the key features that emerge from the competing short- and long-range processes are well visible in 2D, the light-cone propagation of correlations on short times and the supersonic spreading on longer times.
@@
Conclusions: We investigated the dynamics of density-density correlations emerging from the interplay of short- and global-range couplings, finding two key features, a light-cone propagation at short distances and times followed by a supersonic, distance-independent, spreading. The crossover time between these dynamical features and their relative strength depend crucially on the ratio between the global coupling and the short-range tunneling. The general character of our results is highlighted by the robustness to the different quench situations, the dimensionality of the system, the presence of dissipation, or the form of the global coupling. We further checked the stability of the features under local atomic dephasing supp . We identify that the emerging dynamics relies on the presence of fluctuations in the long-range couplings, stemming either from the cavity field or the interaction itself. The role of fluctuations is essential for propagating the correlations outside of the light-cone, which further underline the importance of capturing the fluctuations in the atoms-cavity coupling in obtaining the correct quantum dynamics Halati et al. (2020a, 2022, 2025). We emphasize that our results cannot be inferred from considerations based on general bounds on the propagation of correlations, which would be dominated by the contributions of the global-range couplings, such that numerical simulations of the quantum dynamics are needed. The studied effects could be realized and investigated in current state-of-the-art experiments of ultra-cold atoms coupled to optical cavities. However, due to its generality our findings are applicable to a much wider class of experimental systems. Our results can guide the experimental investigations in controlling the spreading of correlations in platforms with competing interaction scales.
@@
Acknowledgments: We thank J.P. Brantut and T. Donner for fruitful discussions. We acknowledge support by the Swiss National Science Foundation under Division II Grant No. 200020-219400 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277625399-TRR 185 OSCAR (“Open System Control of Atomic and Photonic Matter”, B4), No. 277146847-CRC 1238 (“Control and dynamics of quantum materials”, C05), CRC 1639 NuMeriQS (“Numerical Methods for Dynamics and Structure Formation in Quantum Systems”) – project No. 511713970 , Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”), and under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769. We also acknowledge support within the QuantERA II Programme (project ”QNet: Quantum transport, metastability, and neuromorphic applications in Quantum Networks”), which has received funding from the EU’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733, as well as from the DFG (Project ID 532771420). This research was supported in part by the National Science Foundation under Grants No. NSF PHY-1748958 and PHY-2309135.
@@ A Time-dependent matrix product states method for long-range quantum systems
In order to ensure the convergence of the density-density correlations, $\mathcal{C}_{nn}$, [see main text Eq. (3) for their definition], up to an error of approximately $10^{-5}$ for the times considered in the results presented in the absence of dissipation, we chose the following convergence parameters: a maximal bond dimension of 400400 states, which ensured a truncation error of at most $5\times 10^{-8}$ at the final time, a time-step of $dtJ/\hbar=2.5\times 10^{-3}$, the local Hilbert space of the bosonic atoms is $N_{\text{bos}}=4$, and the adaptive cutoff of the local Hilbert space of the photonic mode ranged between $N_{\text{pho}}=70$ and $N_{\text{pho}}=8$. For the parameters used in the presence of dissipation a truncation error of $10^{-8}$ at the final time was achieved for the mentioned convergence parameters, the presented results are averaged over at least 750 quantum trajectories.
The numerical results for the time-evolution of the atomic model with global range interactions, Eq. (4) in the main text, were obtained with an implementation based on matrix product states of the two-site version of the time-dependent variational principle approach (TDVP) Haegeman et al. (2011, 2016), within the ITensor Library Fishman et al. (2022). The convergence was assured by the following parameters of the method: a maximal bond dimension of 250250 states, which ensured a truncation error of at most $10^{-7}$ at the final time, a time-step of $dtJ/\hbar=5\times 10^{-3}$, and the local Hilbert space of the bosonic atoms are $N_{\text{bos}}=4$.
@@ 1 In the limit of small coupling
$\hat{H}_{\text{int}}=\frac{U}{2}\sum_{j}\hat{n}_{j}(\hat{n}_{j}-1),~\hat{H}_{\text{kin}}=-J\sum_{\langle j,j^{\prime}\rangle}(\hat{b}_{j}^{\dagger}\hat{b}_{j^{\prime}}+\text{H.c.}),$
@@
$\qquad\times e^{-t^{2}/2\hbar[\hat{H}_{\text{kin}}+\hat{H}_{\text{int}},\hat{H}_{\text{ac}}]}\dots,$
@@
$~~\times\left[1+it\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta}\right]e^{-i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}\big\rangle+O(t^{2}\Omega)$
@@
2 Minimal model for the description of the short-time density-density correlation growth
In this section, we aim to capture the minimal ingredients necessary for understanding the distance-independent spreading of the correlations and derive the short-time scaling for the minimal model, $\mathcal{C}_{nn}(d,t)\propto J^{4}\Omega^{4}t^{8}$. As discussed in the main text, when starting from an uncorrelated Fock state the kinetic processes are responsible for creating density-density correlations locally and the coupling to the cavity determines the supersonic spreading. Thus, as a first approximation we neglect the on-site repulsive interactions and the detuning $\delta$ to explain the short time behaviour. In this situation, the 1D version of the Hamiltonian in Eq. (2) of the main text can be written in the momentum basis as
@@ 2 Minimal model for the description of the short-time density-density correlation growth
[…]
@@
where we used Eqs. (B.12) and (B.13) and $^{\sin(2\hat{\theta})=\hbar\hat{\Omega}/\hat{E}}$. Consequently, we can calculate for two even sites $l,m$ the density-density correlations with
@@
[…]
@@ C Treating the cavity mode as a classical field
Figure C2: The space-time propagation of the correlations $\text{Log}[\mathcal{C}_{nn}(d,t)]$, in the presence of dissipation, where in (a) the cavity dynamics has been replaced with a classical field, Eq. (C.1) and Eq. (C.2), and in panel (b) stochastic noise has been added to the classical field evolution, Eq. (C.1) and Eq. (C.3). The panels show the same data as in Fig. C1(c) and Fig. C1(e), but in a logarithmic scale. The blue lines are a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$ and $\Gamma=0$. In the white regions the value of the correlations is smaller than $10^{-4}$. The parameters used are $\hbar\Omega/J=0.99$ $\hbar\Gamma/J=10$, $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$.
@@
The comparison between the exact results and the classical field approach is also shown in Fig. C1. As discussed in the main text, by the coupling to the cavity field for the parameters shown the dynamics of the correlations is dominated by a supersonic propagation [Fig. 3(a) and Fig. LABEL:fig::dissipation_supp(b)], while for the classical field approach we only observe a light-cone spreading of the correlations [Fig. C1(c) and Fig. C1(d)].
@@
where $\xi(t)$ is a random complex number of magnitude 11 sampled from a uniform distribution at each point in time. For the initial state $\langle a\rangle(0)$ we use a uniformly sampled random complex number. We simulate different realizations of the time-dependent noise term, similar to the quantum trajectories approach, and average over the values of the computed observables for the different realizations. We obtain that in the presence of the stochastic noise, Fig. C1(e) and Fig. C1(f) it exhibits for larger distances features of the distance-independent rise of the correlations, as seen in Fig. C1(f) and in the log-scale depiction of the correlations in Fig. C2. This can be understood by the fact that the noise term appearing in Eq. (C.3) translates to a correlated noise term in the evolution of the atoms acting on all sites. However, the contribution from the distance-independent rise of the correlations is much smaller for the shown parameters compared to the full quantum dynamics. This implies that the noise term is insufficient to capture the full dynamics which requires, in particular, also the cavity-mediated transport of density fluctuations.
Figure D1: (a), (c), (e) The propagation of $\mathcal{C}_{nn}(d,t)$, computed with $\hat{H}$, Eq. (B.1), for the atoms-cavity coupling strength $\hbar\Omega/J\in\{0.03,0.42,0.99\}$ and $U/J=16$. The blue line is a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$. (b), (d), (f) The dynamics of $\mathcal{C}_{nn}(d,t)$ for several distances and $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The magenta lines represent algebraically scaling $\propto t^{\alpha}$, with $\alpha\in\{2,4,8\}$. The parameters used are $N=10$ particles, $L=20$ sites, $\hbar\delta/J=2$, $\hbar\Gamma/J=0$.
D Role of on-site interactions
In this section, we show the propagation of correlations in the regime of strong on-site interactions. In Fig. D1 we show the equivalent of Fig. 1 of the main text for $U/J=16$ instead of $U/J=2$. We see the same qualitative behavior as described in the main text, with a light-cone propagation of the correlations for small atoms-cavity coupling [see Fig. D1(a)] followed by a crossover to supersonic, distance-independent, propagation at larger values of the coupling [see Fig. D1(c),(e)]. The value of the atomic on-site interactions does not play a crucial role in the observed phenomenology as they commute with the global-range terms, thus, for the chosen initial state the kinetic terms are the ones which seed the initial correlations, transported afterwards by the global couplings. However, the on-site interactions are important for the details of the dynamics, e.g. the space and time oscillations observed in Fig. D1 and the velocity of the light-cone propagation Cheneau et al. (2012).
E Truncated Wigner approximation (TWA)